# Chain-of-Discussion: A Multi-Model Framework for Complex Evidence-Based Question Answering

### v2

@@ Abstract
Open-ended question answering requires models to find appropriate evidence to form well-reasoned, comprehensive and helpful answers. In practical applications, models also need to engage in extended discussions on potential scenarios closely relevant to the question. With augmentation of retrieval module, open-source Large Language Models (LLMs) can produce coherent answers often with different focuses, but are still sub-optimal in terms of reliable evidence selection and in-depth question analysis. In this paper, we propose a novel Chain-of-Discussion framework to leverage the synergy among multiple open-source LLMs aiming to provide more correct and more comprehensive answers for open-ended QA, although they are not strong enough individually. Our experiments show that discussions among multiple LLMs play a vital role in enhancing the quality of answers. We will release our data and code for further research.
@@ 1 Introduction
Figure 1: The process of Chain-of-Discussion, compared with chain-of-thought. The green parts are necessary to answer user’s question. Blue parts indicate closely related to the question, useful for detailed/extended discussions. Red parts are irrelevant content that should be avoided.
@@ 5.1 Data Collection
To reflect the diversity in practical scenarios, we focus on the fields of marriage, family affairs, and inheritance, covering various legal disputes, e.g., divorce, custody, contracts, property, etc. From a pool of 609 questions collected from real users with consultants’ responses through Web Search Engines, we employ an legal expert to select 200 questions to ensure their semantic distinctiveness and coverage. See more details in Ethics Statement.
@@ 5.2 Experimental Setup
We employ each LLM as the target model, replicating the experimental settings. We report the performance for each LLM as the target role. More details and prompt templates are in Appendix A and  E.
@@
Following previous works Liu et al. (2023); Chan et al. (2024), we employ GPT-4 to evaluate the quality of generated responses, with the expert-written responses, necessary and optional articles as reference. We prompt gpt4-turbo-0125 to score a response in an integer between 1 and 10 based on correctness and comprehensiveness. If there is no clear reason to indicate a response is significantly better or worse than human-written ones, a score of around 7 should be given (Scoring prompt is in Appendix F).
@@ 6.1 Evidence-Centric Evaluation
We accordingly design two metrics of accuracy, N-Acc and O-Acc, to assess the correctness and comprehensiveness, respectively. We utilize the not required articles as negative samples. For N-Acc, we employ the necessary articles as positive samples, while the optional articles for O-Acc. We employ heuristic methods to examine if a response has used an article (details in Appendix C).
@@ 6.3 Limitations of Open-Source LLMs
Reasoning with Knowledge. Similarly, we find the lack of good reasoning capabilities and knowledge may also limit the performance of LLMs in revising evidence analysis. LLMs tend to focus on the literal differences between questions and law articles. Users often omit commonsense information in their questions, whereas law articles exhibit a rigorous and detailed structure. The stylistic differences between questions and evidence may lead to the misjudgments of LLMs (more examples and analysis in Appendix D.2).
@@ Appendix B Details of Evidence Annotation
We assign relevant scores of 2, 1, and 0 to the three categories of articles respectively. If the average score of an article exceeds 1.66, it will be regarded as a necessary one. And the articles with average scores less than 0.67 can be regarded as not required, while the remaining ones are optional.
@@
Appendix C Rules to Identify the Article Reference
@@
Appendix D More Results and Analysis
D.1 Lengths of Generated Responses
@@
D.2 Examples required Reasoning with Commonsense Knowledge
@@ D.3 Significant Test
Appendix E Prompts of Chain-of-Discussion
E.1 Prompt of Question Analysis
To obtain the question analysis, we employ the prompts as below ([]indicates the English translation of the prompts in Chinese):
@@ E.1 Prompt of Question Analysis
咨询者的问题是“[[question]]”
[ The consultant’s question is: “[[question]]”]
下面是检索器提供的参考法条: [[articles]]
[ Below are the reference articles provided by the retriever: [[articles]]]
接下来,请你分析咨询者的问题“[[question]]”请你站在咨询者的角度,首先分析咨询者所处的环境及其面对的客观事实,再围绕着咨询者的问题,分析他可能需要了解哪些方面的法律规定。你需要先对问题进行分析,再分析各个参考法条是否有该问题有关。请你遵循格式,以“问题分析:”、“法条分析:”分别作为两段分析的开头。
[ Next, please analyze the consultant’s question "[[question]]" Please put yourself in the consultant’s shoes, first analyzing the environment they are in and the objective facts they are facing. Then, focusing on the consultant’s question, analyze which aspects of the legal articles they might need to understand. You need to first analyze the question, then assess whether each reference article is relevant to the issue. Please follow the format, starting each section of the analysis with "Question Analysis:" and "Article Analysis:" respectively. ]
@@ E.2 Prompts of Summary
E.3 Prompts of Evidence Analysis
@@
咨询者的问题是“[[question]]”
[ The consultant’s question is:[[question]]]
检索器提供的参考法条: [[articles]]
[ The reference articles provided by the retriever are: [[articles]]]
你需要先对该问题的关键点进行分析,然后再逐个分析每个法条是否对于解答该问题有帮助。分析法条的过程中,请你先思考法条规定了何种权利和义务,或者对何种行为实施了禁止令。如果法条中规定或禁止的内容与问题中的关键点有一定的相关性,那么该法条有可能对于解答问题有所帮助;否则,该法条大概率对解答问题没有帮助。
@@ E.4 Prompt of Critique
问题:[[question]]
[ Question: [[question]]]
法条:[[article]]
[ Articles: [[article]]]
律师对于法条的分析:[[art_ana]]
[ Lawyer’s analysis for the articles:[[art_ana]]]
接下来,请先用简洁的语言点评律师对于[[cur_art_id]]的分析。之后,请你判断他的分析是否误解了法条的内容。
[ Next, please provide a concise critique of the lawyer’s analysis of [[cur_art_id]]. Then, determine whether their analysis misconstrues the content of the legal article. ]
E.5 Prompt of revising
@@
问题:[[question]]
[ Question:[[question]]]
法条:[[article]]
[ Articles[[article]]]
律师对于问题的分析:[[que_ana]]
[ Lawyer’s Analysis of the Question: [[que_ana]]]
律师对于法条的分析:[[art_ana]]
[ Lawyer’s Analysis of the Legal Article: [[art_ana]]]
专家点评:[[critiques]]
[ Expert Critiques: [[critiques]]]
接下来,请你重写一份更为正确的法条分析。在重写后的法条分析的结尾,请你按照你的分析,评估一下该法条是否可能有助于解答问题。
@@ E.6 Prompt of response
UTF8gbsn[ The consultant’s question is: "[[question]]"]
@@
UTF8gbsn[ Below are the reference articles provided by the retriever: [[articles]]]
@@
UTF8gbsn[ Question Analysis: [[que_ana]]]
@@
UTF8gbsn[ Article Analysis: [[art_ana]]]
@@
Appendix F Scoring Prompt of GPT-4
Following CritiqueLLM Ke et al. (2023), we employ a reference-based prompt to instruct GPT-4 to assess the overall quality of the responses generated by open-source LLMs. We use the human-written response and the necessary and option articles as reference. The prompt is shown as below:

### v3

@@ Abstract
Open-ended question answering requires models to find appropriate evidence to form well-reasoned, comprehensive and helpful answers. In practical applications, models also need to engage in extended discussions on potential scenarios closely relevant to the question. With augmentation of retrieval module, open-source Large Language Models (LLMs) can produce coherent answers often with different focuses, but are still sub-optimal in terms of reliable evidence selection and in-depth question analysis. In this paper, we propose a novel Chain-of-Discussion framework to leverage the synergy among multiple open-source LLMs aiming to provide more correct and more comprehensive answers for open-ended QA, although they are not strong enough individually. Our experiments show that discussions among multiple LLMs play a vital role in enhancing the quality of answers.
Chain-of-Discussion: A Multi-Model Framework for Complex Evidence-Based Question Answering
Mingxu Tao1,2,   Dongyan Zhao1,2,3,   Yansong Feng1 ✉ 1Wangxuan Institute of Computer Technology, Peking University 2Center for Data Science, Peking University 3Key Laboratory of Intelligent Press Media Technology, Peking University {thomastao,zhaodongyan,fengyansong}@pku.edu.cn
@@ 1 Introduction
Figure 1: The process of Chain-of-Discussion (bottom part), compared with chain-of-thought (middle part). The green parts are necessary to answer user’s question. Blue parts indicate closely related to the question, useful for detailed/extended discussions. Red parts are irrelevant contents that should be avoided.
@@ 5.1 Data Collection
To reflect the diversity in practical scenarios, we focus on the fields of marriage, family affairs, and inheritance, covering various legal disputes, e.g., divorce, custody, contracts, property, etc. From a pool of 609 questions collected from real users with consultants’ responses through Web Search Engines, we employ an legal expert to select 200 questions [^1] To access the legal consultation data: https://github.com/kobayashikanna01/Chain-of-Discussion to ensure their semantic distinctiveness and coverage. See more details in Ethics Statement.
@@ 5.2 Experimental Setup
We employ each LLM as the target model, replicating the experimental settings. We report the performance for each LLM as the target role. More details and prompt templates are in Appendix A and  F.
@@
Following previous works Liu et al. (2023); Chan et al. (2024), we employ GPT-4 to evaluate the quality of generated responses, with the expert-written responses, necessary and optional articles as reference. We prompt gpt4-turbo-0125 to score a response in an integer between 1 and 10 based on correctness and comprehensiveness. If there is no clear reason to indicate a response is significantly better or worse than human-written ones, a score of around 7 should be given (Scoring prompt is in Appendix G).
@@ 5.3 Main Results
We provide more results for other baseline methods in Appendix C.
@@ 6.1 Evidence-Centric Evaluation
We accordingly design two metrics of accuracy, N-Acc and O-Acc, to assess the correctness and comprehensiveness, respectively. We utilize the not required articles as negative samples. For N-Acc, we employ the necessary articles as positive samples, while the optional articles for O-Acc. We employ heuristic methods to examine if a response has used an article (details in Appendix D).
@@ 6.3 Limitations of Open-Source LLMs
Reasoning with Knowledge. Similarly, we find the lack of good reasoning capabilities and knowledge may also limit the performance of LLMs in revising evidence analysis. LLMs tend to focus on the literal differences between questions and law articles. Users often omit commonsense information in their questions, whereas law articles exhibit a rigorous and detailed structure. The stylistic differences between questions and evidence may lead to the misjudgments of LLMs (more examples and analysis in Appendix E.2).
@@ 7 Conclusions
Acknowledgements
This work is supported in part by NSFC (62161160339) and Beijing Science and Technology Program (Z231100007423011). We thank the anonymous reviewers for their helpful discussions and suggestions. For any correspondence, please contact Yansong Feng.
@@ Appendix B Details of Evidence Annotation
We assign relevant scores of 2, 1, and 0 to the three categories of articles, respectively. If the average score of an article exceeds 1.66, it will be regarded as a necessary one. And the articles with average scores less than 0.67 can be regarded as not required, while the remaining ones are optional.
@@
Appendix C Extended Evaluation
C.1 Results of Other Baselines
In this work, we carefully design a two-stage framework for the legal consultation task, including question analysis and evidence analysis. To give a fair comparison, the chain-of-thought baseline also follows the two-stage settings for generation.
In our early attempt, we also implement a real-vanilla chain-of-thought baseline Wei et al. (2022), which only asks the LLM to directly think whether an article can be used as the evidence and then output the response.
As an ablation study, we also implement a self-criticism baseline Tan et al. (2023). Compared to the mutual-criticism method in Section 4.2, we only ask the target LLM to review whether to revise its evidence analysis, without the help of other LLMs.
Table 5 shows the results of these two baselines. We find the real-vanilla CoT baseline demonstrates much worse performance for all four LLMs. Without the step of question analysis, the 7B LLMs fail to select correct articles, and such failure results in erroneous responses. We can also find a 7B LLM cannot perform well to criticize itself. As we discuss in Section 6.3, the 7B LLMs sometimes fail to provide useful insights about revising its own inaccurate analyses. It is also one of our motivations to introduce the collaboration among multiple models.
Method Baichuan2 Deepseek Qwen Xverse
Vanilla CoT 5.490 6.360 5.545 5.800
Self-Criticism 5.640 6.250 5.790 6.005
Two-Stage CoT 5.750 6.465 5.835 6.015
Our CoD 6.090 6.580 5.955 6.125
Table 5: The average scores of each method with distinct target LLMs (evaluated by GPT-4).
C.2 Results under More Evidences
In the main experiments, we provide five articles for each query, including 2.75 required articles on average. We examine whether Chain-of-Discussion can be effective when provided more candidate articles. Thus, we add extra articles, changing the number of given evidences to 7 or 10.
Table 6 shows the results when we provide 5, 7 or 10 articles as evidence, respectively. As the number of articles increases, the qualities of the responses generated by both baseline and Chain-of-Discussion become a bit worse. It may be attributed to the noise brought by more irrelevant articles, which results in more incorrect contents in model’s responses.
We also find when the number of given articles increases, Chain-of-Discussion can bring more improvement to both Baichuan2-7B and Deepseek-7B, which are the worst-performing and the best-performing models in this work. We believe that Chain-of-Discussion can better filter irrelevant evidence from the given articles.
Num. of Articles 5 7 10
Baichuan2+CoT 5.750 5.680 5.575
Baichuan2+CoD 6.090 6.075 6.030
$\Delta(\text{CoD}-\text{CoT})$ +0.280 +0.395 +0.455
Deepseek+CoT 6.465 6.435 6.380
Deepseek+CoD 6.580 6.560 6.530
$\Delta(\text{CoD}-\text{CoT})$ +0.115 +0.125 +0.150
Table 6: The average scores of Baichuan2-7B and Deepseek-7B with different numbers of candidate articles (evaluated by GPT-4).
Appendix D Rules to Identify the Article Reference
@@
Appendix E More Analysis
E.1 Lengths of Generated Responses
@@
E.2 Examples required Reasoning with Commonsense Knowledge
@@ E.3 Significant Test
Appendix F Prompts of Chain-of-Discussion
F.1 Prompt of Question Analysis
To obtain the question analysis, we employ the prompts as below ([__text__] indicates the English translation of the prompts in Chinese):
@@ F.1 Prompt of Question Analysis
{CJK*}
UTF8gbsn咨询者的问题是“{{question}}”
[ The consultant’s question is: “{{question}}”]
{CJK*}
UTF8gbsn下面是检索器提供的参考法条: {{articles}}
[ Below are the reference articles provided by the retriever: {{articles}} ]
{CJK*}
UTF8gbsn接下来,请你分析咨询者的问题“{{question}}”请你站在咨询者的角度,首先分析咨询者所处的环境及其面对的客观事实,再围绕着咨询者的问题,分析他可能需要了解哪些方面的法律规定。你需要先对问题进行分析,再分析各个参考法条是否有该问题有关。请你遵循格式,以“问题分析:”、“法条分析:”分别作为两段分析的开头。
[ Next, please analyze the consultant’s question "{{question}}" Please put yourself in the consultant’s shoes, first analyzing the environment they are in and the objective facts they are facing. Then, focusing on the consultant’s question, analyze which aspects of the legal articles they might need to understand. You need to first analyze the question, then assess whether each reference article is relevant to the issue. Please follow the format, starting each section of the analysis with "Question Analysis:" and "Article Analysis:" respectively. ]
@@ F.2 Prompt of Summary
F.3 Prompt of Evidence Analysis
@@
{CJK*}
UTF8gbsn咨询者的问题是“{{question}}”
[ The consultant’s question is: {{question}} ]
{CJK*}
UTF8gbsn检索器提供的参考法条: {{articles}}
[ The reference articles provided by the retriever are: {{articles}} ]
{CJK*}
UTF8gbsn你需要先对该问题的关键点进行分析,然后再逐个分析每个法条是否对于解答该问题有帮助。分析法条的过程中,请你先思考法条规定了何种权利和义务,或者对何种行为实施了禁止令。如果法条中规定或禁止的内容与问题中的关键点有一定的相关性,那么该法条有可能对于解答问题有所帮助;否则,该法条大概率对解答问题没有帮助。
@@ F.4 Prompt of Critique
{CJK*}
UTF8gbsn问题:{{question}}
[ Question: {{question}} ]
{CJK*}
UTF8gbsn法条:{{article}}
[ Articles: {{article}} ]
{CJK*}
UTF8gbsn律师对于法条的分析:{{art_ana}}
[ Lawyer’s analysis for the articles: {{art_ana}} ]
{CJK*}
UTF8gbsn接下来,请先用简洁的语言点评律师对于{{cur_art_id}}的分析。之后,请你判断他的分析是否误解了法条的内容。
[ Next, please provide a concise critique of the lawyer’s analysis of Article No. {{cur_art_id}}. Then, determine whether their analysis misconstrues the content of the legal article. ]
F.5 Prompt of Revising
@@
{CJK*}
UTF8gbsn问题:{{question}}
[ Question: {{question}} ]
{CJK*}
UTF8gbsn法条:{{article}}
[ Articless: {{article}} ]
{CJK*}
UTF8gbsn律师对于问题的分析:{{que_ana}}
[ Lawyer’s Analysis of the Question: {{que_ana}} ]
{CJK*}
UTF8gbsn师对于法条的分析:{{art_ana}}
[ Lawyer’s Analysis of the Legal Article: {{art_ana}} ]
{CJK*}
UTF8gbsn专家点评:{{critiques}}
[ Expert Critiques: {{critiques}} ]
{CJK*}
UTF8gbsn接下来,请你重写一份更为正确的法条分析。在重写后的法条分析的结尾,请你按照你的分析,评估一下该法条是否可能有助于解答问题。
@@ F.6 Prompt of Response
UTF8gbsn[ The consultant’s question is: {{question}} ]
@@
UTF8gbsn[ Below are the reference articles provided by the retriever: {{articles}} ]
@@
UTF8gbsn[ Question Analysis: {{que_ana}} ]
@@
UTF8gbsn[ Article Analysis: {{art_ana}} ]
@@
Appendix G Scoring Prompt of GPT-4
Following CritiqueLLM Ke et al. (2023), we employ a reference-based prompt to instruct GPT-4 to assess the overall quality of the responses generated by open-source LLMs. We use the human-written response and the necessary and option articles as reference for evaluation. The prompt is shown as below: